# Not All Language Model Features Are Linear

### v2

@@ 5.1 Intervening on Circular Day and Month Representations
Figure 5: Visual representation of the intervention process.
Figure 6: Mean and 96% error bars for intervention methods.
@@
2. Intervene on the subspace: Say our initial prompt had $\alpha=\alpha_{j}$ and we are intervening with $\alpha=\alpha_{j^{\prime}}$. In this step, we replace the model’s projection on the subspace ${\mathbf{P}}\cdot{\mathbf{W}}_{i,l}$, which will be close to $\texttt{circle}(\alpha_{j}$), with the “clean” point $\texttt{circle}(\alpha_{j^{\prime}})$. Note that we do not use the hidden state $\mathbf{x}^{j^{\prime}}_{i,l}$ from the “clean” run, only the “clean” label $\alpha_{j^{\prime}}$. In practice, other subspaces of $\mathbf{x}^{j}_{i,l}$ are used by the model in alternate pathways to compute the answer, so if we just intervene on the circular subspace the logit difference effect is not as pronounced. To solve this problem, we average out the portion of the activation not in the intervened subspace. Letting $\overline{{\bf x}_{i,l}}$ be the average of ${\bf x}^{j}_{i,l}$ across all prompts indexed by $j$ and ${\mathbf{P}}^{+}$ be the pseudoinverse of ${\mathbf{P}}$, we intervene via the formula
@@
We run our patching on all $49$ Weekday problems and $144$ Month problems and use as “clean” runs the $6$ or $11$ other possible values for $\beta$, resulting in a total of $49*6$ patching experiments for Weekdays and $144*11$ patching experiments for Months. We also run baselines where we
Figure 7: Off distribution interventions on Mistral layer $5$ on the Weekdays task. The color corresponds to the highest logit $\gamma$ after performing our circular subspace intervention process on that point.
(1) replace the entire subspace corresponding to the first $5$ PCA dimensions with the corresponding subspace from the clean run, (2) replace the entire layer with the corresponding layer from the clean run, and (3) replace the entire layer with the average across the task. The metric we use is average logit difference across all patching experiments between the original correct token ($\alpha_{j}$) and the target token ($\alpha_{j^{\prime}}$). See Fig. 6 for these interventions on all layers of Mistral 7B and Llama 3 8B on Weekdays and Months.
The main takeaway from Fig. 6 is that circular subspaces are causally implicated in computing $\gamma$, especially for Weekdays. Across all models and tasks, early layer interventions on the circular subspace have almost the same intervention effect as patching the entire layer, and are sometimes even better than patching the top PCA dimensions from the clean problem. Note that patching experiments in Appendix I show $\alpha$ is copied to the final token on layers $15$ to $17$, which is why interventions drop off there.
@@
5.2 Uncovering Output Representations Using Regression
So far, we have focused on examining and intervening on the representation for $\alpha$, which was conveniently present as a circle in the top PCA components on top of the $\alpha$ token. In this section, we will examine how the generated output, $\gamma$, is represented.
First, to isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Appendix J, show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to what Stolfo et al. (2023) find in prior work studying arithmetic circuits). Thus, we know where to look for a representation of $\gamma$: in the second half of the layers on the token before $\gamma$. However, patching alone cannot tell use how $\gamma$ is represented.
More details and experiments using EVR to completely break down language model states are in Appendix J. Here, we use EVR to determine the representation for $\gamma$ by plotting the top two PCA components of the layer $25$ Mistral 7B activations after subtracting the components that can be explained using a regression with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$). The result, shown in Fig. 8, is an incredibly clear circle in $\gamma$, which suggests that the model’s generated representation of $\gamma$ lies along a circle. A simple PCA projection was not enough to find this result because the representation for $\gamma$ has interference from $\alpha$ and $\beta$, which the EVR removes. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” Nanda et al. (2023a) or “pizza” Zhong et al. (2024) algorithm in late MLP layers.
@@ 5.3 Continuity of Circular Representations
In past sections, the representations of the interpretable numeric quantities we have discovered have been mostly discontinuous; that is, the days of the week and months of the year in Fig. 1 and Fig. 15 are clustered at the vertices of a heptagon and dodecagon, and there is nothing "between" adjacent weekdays or months along the circle. Although continuity of the representation is not a requirement of Definition 3, it would further decrease the $\epsilon$-mixture index, and would also increase our subjective perception of the circular feature as an intrinsic model feature representing a continuous quantity (time). Thus, we create a synthetic dataset containing the text "[very early/very late] on [Monday/Tuesday/…/Sunday]" and simply plot the projections of the layer $30$ activations into the top two PCA components of the activations of [Monday/Tuesday/…/Sunday]. The results, shown in Fig. 9, show that Mistral 7B indeed can map intermediate quantities to their expected place in the circle: the very early and very late version of each weekday are more towards the last and the next weekday along the circle, respectively. We show similar results for "[morning/evening] on [Monday/Tuesday/…/Sunday]" in Appendix Fig. 19.
@@ 6 Discussion
It is unclear why we did not find more interpretable multi-dimensional features. We are unsure if we are failing to interpret some of the high-scoring multi-dimensional features, if our clustering technique is not powerful enough to find some features, or if there are truly not that many. Our definition for an irreducible feature (Definition 2) also had to be relaxed to hold in practice (Definition 3). Thus, although this work provides preliminary evidence for the multi-dimensional superposition hypothesis (2), it is still unclear if this theory provides the best description for the representations models use. We also did not find a small subset of MLP neurons implementing the “clock” algorithm, leaving as an open question to what extent models use multi-dimensional representations for algorithmic tasks. Finally, we only ran experiments on models up to size 8B; however, recent work (Huh et al., 2024) implies representations may become universal with growing model size.
@@ B.2.2 $\epsilon$-Mixture Index
Figure 10: Testing irreducibility of synthetic features. Left in each subfigure: Distributions of $\mathbf{x}$. For feature ${\mathbf{a}}$, 36.21% lies within the narrow dotted lines, indicating the feature is likely a mixture. For feature ${\mathbf{b}}$, 17.84% lies within the wide lines, indicating the feature is unlikely to be a mixture. The green cross indicates the angle $\theta$ that minimizes mutual information. Middle in each subfigure: Histograms of the distribution of $\mathbf{v}\cdot\mathbf{x}$ with red lines indicating a $2\epsilon$-wide region. Right in each subfigure: Mutual information between $\mathbf{a}$ and $\mathbf{b}$ as a function of the rotation angle $\theta$ of matrix $\mathbf{R}$. Both features have a large ($\geq 0.5$ bits) minimum mutual information and so are likely not separable.
@@ Appendix D Toy Case of Training SAEs on Circles
First however, consider activations for a single circle – points uniformly distributed on the unit circle in $\mathbb{R}^{2}$. We train SAEs on this data with encoder $\texttt{Enc}(\mathbf{x})=\texttt{ReLU}(\mathbf{W_{e}}(\mathbf{x}-\mathbf{b_{d}})+\mathbf{b_{e}})$ and decoder $\texttt{Dec}(\mathbf{f})=\mathbf{W_{d}}\mathbf{f}+\mathbf{b_{d}}$. We train SAEs with $m=2$ and $m=10$ with the Adam optimizer and a learning rate of $10^{-3}$, sparsity penalty $\lambda=0.1$, for 20,000 steps, and a warmup of 1000 steps. In Fig. 12 we show the dictionary elements of these SAEs. When $m=2$, the network must use both SAE features on each input point, and uses $\mathbf{d_{b}}$ to shift the reconstructed circle so it is centered at the origin. When $m=10$, $\mathbf{d_{b}}\approx 0$ and the features spread out across the circle having close neighbors, with only a subset being active on any one input.
Figure 12: SAEs trained to reconstruct a single 2d circle with $m=2$ (left) and $m=10$ (middle and right) dictionary elements. When there are several SAE features, there is not a natural choice feature directions, and the dictionary elements spread out across the circle.
@@ F.2 Mistral 7B methods and results
Appendix G Further Experiment Details
G.1 Assets Information
@@ G.3 Error Bar Calculation
Appendix H More Weekdays and Months Plots and Details
@@ Appendix I Patching
To isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Fig. 18 show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to prior work studying arithmetic circuits (Stolfo et al., 2023)). Moreover, fine-grained patching in Appendix J shows that there are just a few responsible attention heads for the writes to the token before $\gamma$. However, patching alone cannot tell use how or where $\gamma$ is represented. For that, we need a new technique, which we expand on in the next section.
Appendix J Explanation via Regression (EVR)
In this section, we introduce a new technique for empirically explaining hidden representations in algorithmic problems: explanation via regression (EVR). Given a set of tokens $T$ with a corresponding set of hidden states $X_{i,l}$, we explain the variance in $X_{i,l}$ by adding together hand-chosen functions of $t$. This gives us an explanation of what the transformation $T\rightarrow X_{i,l}$ computes. For a given choice of explanation functions $\{{\mathbf{g}}_{i}(t)\}$, the $r^{2}$ value of a linear regression from $\{{\mathbf{g}}_{i}(t)\}$ to $X_{i,l}$ gives a measure of the explained variance in $X_{i,l}$. But what functions ${\mathbf{g}}_{i}$ should we choose?
@@
We build a list of ${\mathbf{g}}_{i}$ iteratively and greedily. At each iteration, we perform a linear regression with the current list ${\mathbf{g}}_{1}\dots\mathbf{g}_{k}$, visualize and interpret the residual prediction errors, and build a new function ${\mathbf{g}}_{k+1}$ representing these errors to add to the list. Once most variance is explained, we can conclude that ${\mathbf{g}}_{1},\dots,{\mathbf{g}}_{k}$ constitutes the entirety of what is represented in the hidden states. This information tells us what can and cannot be extracted via a linear probe, without having to train any probes. Furthermore, if we treat each ${\mathbf{g}}_{i}$ as a feature (see Definition 1), then the linear regression coefficients tell us which directions in $X_{i,l}$ these features are represented in, connecting back to 2.
We apply EVR to Months and Weekdays. Since $X_{i,l}$ consists of modular addition problems with two inputs $\alpha$ and $\beta$, we can visualize the errors as we iteratively construct ${\mathbf{g}}_{1},\ldots,{\mathbf{g}}_{k}$ by making a heatmap with $\alpha$ and $\beta$ on the two axes, where the color shows what kind of error is made. More specifically, we take the top 3 PCA components of the error distribution and assign them to the colors red, green, and blue. We call the resulting heatmap a residual RGB plot. Errors that depend primarily on $\alpha$, $\beta$, or $\gamma$ show up as horizontal, vertical, or diagonal stripes on the residual RGB plot.
In  Fig. 20, we perform EVR on the layer 17-29 hidden states of Mistral 7B on the Weekdays task; additional deconstructions are in Appendix J. We find that a circle in $\gamma$ develops and grows in explanatory power; we plot the layer $25$ residuals after explaining with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$) in Fig. 8 to show this incredibly clear circle in $\gamma$. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” (Nanda et al., 2023a) or “pizza” (Zhong et al., 2024) algorithm in late MLP layers.

### v3

@@ 5.1 Intervening on Circular Day and Month Representations
Figure 5: Visual representation of the circular intervention process. Top: We learn a circular probe on the PCA projection of a training set. Bot: To intervene, we change the circular representation to $\alpha_{j}^{\prime}$ and average ablate other dimensions.
Figure 6: Mean and 96% error bars for intervening on the $\alpha$ token across layers using different intervention methods. The circular intervention technique outperforms patching only the top $5$ PCA components and leaving the rest unchanged, and almost reaches the upper bound performance of patching the entire layer.
@@
2. Intervene on the subspace: Say our initial prompt had $\alpha=\alpha_{j}$ and we are intervening with $\alpha=\alpha_{j^{\prime}}$. In this step, we replace the model’s projection on the subspace ${\mathbf{P}}\cdot{\mathbf{W}}_{i,l}$, which will be close to $\texttt{circle}(\alpha_{j}$), with the “clean” point $\texttt{circle}(\alpha_{j^{\prime}})$. Note that we do not use the hidden state $\mathbf{x}^{j^{\prime}}_{i,l}$ from the “clean” run, only the “clean” label $\alpha_{j^{\prime}}$. In practice, other subspaces of $\mathbf{x}^{j}_{i,l}$ may be used concurrently by the model in “backup” circuits (see e.g. Wang et al. (2022)) to compute the answer, so if we just intervene on the circular subspace the remaining components of the activation may interfere in downstream computations. Thus, to isolate the effect of our intervention, we set the average ablate the portion of the activation not in the intervened subspace. Letting $\overline{{\bf x}_{i,l}}$ be the average of ${\bf x}^{j}_{i,l}$ across all prompts indexed by $j$ and ${\mathbf{P}}^{+}$ be the pseudoinverse of ${\mathbf{P}}$, we intervene via the formula
@@
We run our patching on all $49$ Weekday problems and $144$ Month problems and use as “clean” runs the $6$ or $11$ other possible values for $\beta$, resulting in a total of $49*6$ patching experiments for Weekdays and $144*11$ patching experiments for Months. We also run baselines where we (1) replace the entire subspace corresponding to the first $5$ PCA dimensions with the corresponding subspace from the clean run, (2) replace the entire layer with the corresponding layer from the clean run, and (3) replace the entire layer with the average across the task. The metric we use is average logit difference across all patching experiments between the original correct token ($\alpha_{j}$) and the target token ($\alpha_{j^{\prime}}$). See Fig. 6 for these interventions on all layers of Mistral 7B and Llama 3 8B on Weekdays and Months.
Figure 7: Off distribution interventions on Mistral layer $5$ on the Weekdays task. The color corresponds to the highest logit $\gamma$ after performing the circular subspace intervention on that point.
The main takeaway from Fig. 6 is that circular subspaces are causally implicated in computing $\gamma$, especially for Weekdays. Across all models and tasks, early layer interventions on the circular subspace have almost the same intervention effect as patching the entire layer, and are usually better than patching the top PCA dimensions from the clean problem.
Patching experiments in Appendix J show $\alpha$ is copied to the final token on layers $15$ to $17$, which is why interventions drop off there. Additionally, while in this section we train a probe on a dataset of prompts, in Section 5.2, we show that intervening on the circle discovered via SAE clustering in Section 4 also works.
@@
5.2 Intervening with the SAE Plane
In the last section, we train probes manually on the PCA of the activations to fit a circle. A perhaps more natural approach is intervening with the precise circle we found in Section 4. To determine if this approach is feasible, we first project layer 8 Mistral 7B Weekdays activations into the weekdays plane that was discovered by clustering (see Fig. 15; the plane is defined by PCA dimensions 2 and 3 of the cluster). In the rest of this section, we call this plane the SAE plane. In Fig. 19, we find that indeed, the Weekdays representations projected into the SAE plane form a circle (see Fig. 19). We thus can fit a circular probe to this 2D plane as in Eq. 5. Similarly, we call this probe the SAE probe.
Figure 8: Interventions on the Mistral 7B Weekdays task with different methods of determining the probe.
Because we only have layer 8 clustering results for Mistral, we train an SAE probe only on layer 8. We then evaluate interventions using this SAE probe on layer 8 of Mistral, but also at neighboring layers, since nearby layers should have similar representations (see e.g. Belrose et al. (2023)). We compare to two baselines: 1) training a normal circular probe on the PCA projections of each layer as described in Section 5.1, and 2) training a circular probe only on layer 8 and then evaluating on adjacent layers (in the same way as for the layer 8 SAE probe). We show the results of these methods in Fig. 8.
We find that on all layers, using the SAE probe only slightly decreases intervention performance as compared to training a circular probe (from -2.58 to -2.01 average logit difference on layer 8). Even more interestingly, the layer 8 SAE probe is much more robust to layer shifts than the layer 8 circular probe; for example, using the layer 8 circular probe on layer 6 results in an average logit difference of 0.029, whereas using the layer 8 SAE probe results in an average logit difference of -2.32. This is intriguing evidence that the SAE is perhaps finding more “true” (or at least more robust) features than our circular probing technique.
@@ 5.3 Continuity of Circular Representations
In past sections, the representations of the interpretable numeric quantities we have discovered have been mostly discontinuous; that is, the days of the week and months of the year in Fig. 1 and Fig. 15 are clustered at the vertices of a heptagon and dodecagon, and there is nothing "between" adjacent weekdays or months along the circle. In this section, we will examine the "continuity" of the circular features we have discovered. Although continuity of the representation is not a requirement of Definition 3, it would further decrease the $\epsilon$-mixture index, and would also increase our subjective perception of the circular feature as an intrinsic model feature representing a continuous quantity (time). Thus, we create a synthetic dataset containing the text "[very early/very late] on [Monday/Tuesday/…/Sunday]" and simply plot the projections of the layer $30$ activations into the top two PCA components of the activations of [Monday/Tuesday/…/Sunday]. The results, shown in Fig. 9, show that Mistral 7B indeed can map intermediate quantities to their expected place in the circle: the very early and very late version of each weekday are more towards the last and the next weekday along the circle, respectively. We show similar results for "[morning/evening] on [Monday/Tuesday/…/Sunday]" in Appendix Fig. 22.
@@ 6 Discussion
It is unclear why we did not find more interpretable multi-dimensional features. We are unsure if we are failing to interpret some of the high-scoring multi-dimensional features, if most multi-dimensional features lie in dimensions higher than two, if our clustering technique is not powerful enough to find some features, or if there are truly not that many. Additionally, our definitions for irreducible features (Definition 2) are purely statistical and not intervention based, and also had to be relaxed to hold in practice, resulting in measures that return a possibly subjective “degree” of reducibility (Definition 3). Thus, although this work provides preliminary evidence for the multi-dimensional superposition hypothesis (2), it is still unclear if this theory provides the best description for the representations models use. Future work might make progress on this question by investigating new techniques for decomposing model representations, exploring higher dimensional representations, or determining conclusively whether models use representations in ways that necessitate the representations are non-linear.
@@ B.2.2 $\epsilon$-Mixture Index
Figure 10: Testing irreducibility of synthetic features. Left in each subfigure: Distributions of $\mathbf{x}$. For feature ${\mathbf{a}}$, 63.96% lies within the narrow dotted lines, indicating the feature is likely a mixture. For feature ${\mathbf{b}}$, 17.84% lies within the wide lines, indicating the feature is unlikely to be a mixture. The green cross indicates the angle $\theta$ that minimizes mutual information. Middle in each subfigure: Histograms of the distribution of $\mathbf{v}\cdot\mathbf{x}$ with red lines indicating a $2\epsilon$-wide region. Right in each subfigure: Mutual information between $\mathbf{a}$ and $\mathbf{b}$ as a function of the rotation angle $\theta$ of matrix $\mathbf{R}$. Feature ${\mathbf{b}}$ has a large minimum mutual information so is unlikely to be separable; feature ${\mathbf{a}}$ has a medium value of minimum mutual information of about $0.37$ bits.
@@ Appendix D Toy Case of Training SAEs on Circles
First however, consider activations for a single circle – points uniformly distributed on the unit circle in $\mathbb{R}^{2}$. We train SAEs on this data with encoder $\texttt{Enc}(\mathbf{x})=\texttt{ReLU}(\mathbf{W_{e}}(\mathbf{x}-\mathbf{b_{d}})+\mathbf{b_{e}})$ and decoder $\texttt{Dec}(\mathbf{f})=\mathbf{W_{d}}\mathbf{f}+\mathbf{b_{d}}$. We train SAEs with $m=2$ and $m=10$ with the Adam optimizer and a learning rate of $10^{-3}$, sparsity penalty $\lambda=0.1$, for 20,000 steps, and a warmup of 1000 steps. In Fig. 12 we show the dictionary elements of these SAEs. When $m=2$, the SAE must use both SAE features on each input point, and uses $\mathbf{d_{b}}$ to shift the reconstructed circle so it is centered at the origin. When $m=10$, the SAE learns $\mathbf{d_{b}}\approx 0$ and the features spread out across the circle, arranged close together, and only a subset are active on each input.
Figure 12: SAEs trained to reconstruct a single 2d circle with $m=2$ (left) and $m=10$ (middle and right) dictionary elements. When there are several SAE features, there is no natural/canonical choice of feature directions, and the dictionary elements spread out across the circle.
@@ Appendix F GPT-2 and Mistral 7B Dictionary Element Clustering
In this section, we first present pseudocode in Algorithm 1 for the overall high level technique that finds multi-dimensional features and that uses clustering as a subroutine. We then provide the specific clustering algorithm implementations we use for GPT-2 and Mistral.
Input: Dictionary elements $D$, activation vectors $X_{i,l}$, SAE
Output: Irreducible multi-dimensional features
$S_{i,j}\leftarrow\textnormal{{CosineSim}}(D_{i},D_{j})$;
$clusters\leftarrow$ Cluster($S$);
$reconstructions\leftarrow\{\}$;
for $cluster$ in $clusters$ do
       $R_{cluster}\leftarrow$ ids of dictionary elements in $cluster$;
       for ${\bm{x}}_{i,l}$ in $X_{i,l}$ do
             $encoding\leftarrow\texttt{ReLU}(E\cdot{\bm{x}}_{i,l})$;
             if $\max(encoding[R_{cluster}])>0$ then
                   $r\leftarrow D[:,R_{cluster}]\cdot encoding$;
                   $reconstructions\leftarrow reconstructions\cup\{r\}$;
             end if
       end for
end for
$features\leftarrow\{\}$;
for $R$ in $reconstructions$ do
       $proj\leftarrow$ PCA($R$);
       if TestIrreducible($proj$) then
             Add $proj$ to $features$;
       end if
end for
return $features$
Algorithm 1 High Level Clustering Approach For Finding Multi-D Features
@@ F.2 Mistral 7B methods and results
We also investigate the sensitivity of this method to $\tau$ and $k$ by varying $\tau$ and $k$ and showing the max Jaccard similarity between any of the resulting clusters and the days of the week cluster we show in Fig. 15. We show the results in Fig. 16, where we find that varying $k$ has minimal effect, while varying $\tau$ shows $3$ regimes: small $\tau$ causes all features to group in one cluster, so the days of the week cluster is not found; medium $\tau$ causes the days of the week cluster to become identifiable; large $\tau$ causes all features to be divided into their own clusters.
@@
Figure 16: Hyperparameter regimes where the days of the week cluster exists. The cluster exists in the regime between all features clumping together and all features being in their own cluster; this regime seems reasonably stable.
@@
Appendix G Other discovered clusters
In Fig. 17, we plot the top $11$ ranked clusters by the product of a) the measured separability index and b) one minus the measured $\epsilon$-mixture index with $\epsilon=0.1$ (this is just one of many possible ways to get an ordered ranking from a two-parameter score). We color by both the current token (which results in clear patterns for all tokens) and the next token (to see if we find belief states as found by Shai et al. (2024) in toy transformers). We note that weekdays are ranked 9 and so are shown in the plot. Additionally, the next token patterns of the ‘such’ cluster and the ‘B’ cluster do seem to display some clustering independently of the the current token pattern, which might lend the belief state hypothesis some support.
Appendix H Further Experiment Details
H.1 Assets Information
@@ H.3 Error Bar Calculation
Appendix I More Weekdays and Months Plots and Details
Figure 19: Projections of Mistral 7B Weekdays task activations at layer 8 into the plane discovered by clustering layer 8 SAE features.
I.0.1 Basic Plots
I.0.2 Intervening with the SAE Probe
We show the results of projecting Mistral Weekdays representations into the plane discovered by clustering SAE features in Fig. 19. The result is clearly circular.
@@ I.0.2 Intervening with the SAE Probe
I.0.3 Basic Patching
@@ I.0.4 Circle Continuity
To isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Fig. 21 show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to prior work studying arithmetic circuits (Stolfo et al., 2023)). Moreover, fine-grained patching in Appendix K shows that there are just a few responsible attention heads for the writes to the token before $\gamma$. However, patching alone cannot tell use how or where $\gamma$ is represented. For that, we need a new technique, which we expand on in the next section.
Appendix K Explanation via Regression (EVR)
So far, we have focused on examining and intervening on the representation for $\alpha$, which we present as a circle in the top PCA components on top of the $\alpha$ token. In this section, we examine how the generated output, $\gamma$, is represented.
First, to isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Fig. 21 and Fig. 20, show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to what Stolfo et al. (2023) find in prior work studying arithmetic circuits). Thus, we know where to look for a representation of $\gamma$: in the second half of the layers on the token before $\gamma$. However, patching alone cannot tell use how $\gamma$ is represented.
K.1 Using EVR to uncover a circular representation for $\gamma$
We first use EVR to determine the representation for $\gamma$ by plotting the top two PCA components of the layer $25$ Mistral 7B activations after subtracting the components that can be explained using a regression with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$). The result, shown in Fig. 23, is an incredibly clear circle in $\gamma$, which suggests that the model’s generated representation of $\gamma$ lies along a circle. A simple PCA projection was not enough to find this result because the representation for $\gamma$ has interference from $\alpha$ and $\beta$, which the EVR removes. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” Nanda et al. (2023a) or “pizza” Zhong et al. (2024) algorithm in late MLP layers.
K.2 More Experiments with EVR
@@
We now apply EVR to Months and Weekdays to break down $X_{i,l}$ completely into interpretable functions. We build a list of ${\mathbf{g}}_{i}$ iteratively and greedily. At each iteration, we perform a linear regression with the current list ${\mathbf{g}}_{1}\dots\mathbf{g}_{k}$, visualize and interpret the residual prediction errors, and build a new function ${\mathbf{g}}_{k+1}$ representing these errors to add to the list. Once most variance is explained, we can conclude that ${\mathbf{g}}_{1},\dots,{\mathbf{g}}_{k}$ constitutes the entirety of what is represented in the hidden states. This information tells us what can and cannot be extracted via a linear probe, without having to train any probes. Furthermore, if we treat each ${\mathbf{g}}_{i}$ as a feature (see Definition 1), then the linear regression coefficients tell us which directions in $X_{i,l}$ these features are represented in, connecting back to 2.
Since $X_{i,l}$ consists of modular addition problems with two inputs $\alpha$ and $\beta$, we can visualize the errors as we iteratively construct ${\mathbf{g}}_{1},\ldots,{\mathbf{g}}_{k}$ by making a heatmap with $\alpha$ and $\beta$ on the two axes, where the color shows what kind of error is made. More specifically, we take the top 3 PCA components of the error distribution and assign them to the colors red, green, and blue. We call the resulting heatmap a residual RGB plot. Errors that depend primarily on $\alpha$, $\beta$, or $\gamma$ show up as horizontal, vertical, or diagonal stripes on the residual RGB plot.
In  Fig. 24, we perform EVR on the layer 17-29 hidden states of Mistral 7B on the Weekdays task; additional deconstructions are in Appendix K. We find that a circle in $\gamma$ develops and grows in explanatory power; we plot the layer $25$ residuals after explaining with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$) in Fig. 23 to show this incredibly clear circle in $\gamma$. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” (Nanda et al., 2023a) or “pizza” (Zhong et al., 2024) algorithm in late MLP layers.